OpenStudy (anonymous):

A ray of yellow light from a sodium discharge lamp falls on a surface of a diamond in air at incident angle 50°. If at that frequency, 5.24 *1014 Hz, n(diamond) = 2.4, what is the angular deviation of the ray, in degrees, suffered upon transmission? Answer

OpenStudy (noelgreco):

Use Snell's Law.

OpenStudy (anonymous):

but why would they have given me frequency?

OpenStudy (anonymous):

the index of refraction of a material is not the same for every wavelength or frequency so they are saying that for that particular frequency (yellow light) is 2.4. Make sense?

OpenStudy (kyosuke):

Mira, sabemos que la frecuencia del rayo se mantiene constate ella "no depende del medio" por tanto si lo cambias de un medio a otro se mantendrà constante por eso no es tan relevante, personalmente diría que la frecuencia es un distracción del problema que te están dando, calcula mediante la ley de Snell (1)*sin(50)=2.4*sin(x) con eso te debería de salir la respuesta. suerte!.

OpenStudy (anonymous):

@azolotor Thank you for the response, but I cannot find anywhere an equation that relates snell's law to the frequency of light :/

OpenStudy (anonymous):

No no they give you the index of refraction. You can use that directly in Snell's law. I was just explaining where they got their n from and why they told you it was related to its frequency. Frequency is not related to snell's law it is just related to the index of refraction

OpenStudy (anonymous):

Ohhh, ok. So I am correct in the following relation that kyosuke has already pointed out:(1)*sin(50)=2.4*sin(x), where x is the transmission angle? I got 18 degrees.

OpenStudy (anonymous):

yup

OpenStudy (kyosuke):

Como te digo, la frecuencia se mantiene constante lo que varia es la velocidad de la onda electromagnética ( la luz amarilla). De hecho tu solo has visto la ley de Snell de forma simplificada si te das cuenta en indice de refracción se calcula de esta forma n=C/V donde C: velocidad de la luz en el vacio V: velocidad de la luz en el medio Ahora si recuerdas la V de una onda se calcula de la siguiente forma V=(longitud de onda)/f " y ahí esta la frecuencia" pero como es constante se simplificara con la frecuencia de la velocidad de la luz (ya que es la misma) *cualquier duda házmela saber Suerte!.

OpenStudy (anonymous):

thank you @kyosuke @azolotor and @NoelGreco !!! I am just wanting to get all the points I can for my optics class! you guys all helped out so much!!